Circular No. 8412

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 S1 (VAN NESS) M. E. Van Ness, Lowell Observatory, reports his discovery of a comet

on CCD images taken by himself with the 0.59-m Schmidt LONEOS telescope (discovery observation below), reporting an 8′′ moderately condensed coma with a fan-shaped tail 155′′ long in p.a. 225^{o}. B. A. Skiff reports that LONEOS images taken by himself on Sept. 27.4 UT show a nearly circular, poorly condensed coma of red mag 18.0 and ∼ 12′′ across in bright moonlight. Following posting on “The NEO Confirmation Page”, J. E. McGaha (Tucson, AZ, 0.36-m f/10.0 Schmidt-Cassegrain reflector + CCD) reports that 20 stacked 60-s exposures from Sept. 27.3 show an unusual appearance, with a split fan-shaped tail 18′′ long in p.a. 280^{o} and 250^{o}, and an elongated coma in p.a. 260^{o} and no sharp central condensation; McGaha’s 30 stacked exposures on Sept. 28.3 also show a diffuse, elongated coma 14′′ (axis along p.a. 265^{o}–85^{o}) × 4′′, again with no sharp condensation, and an 8′′ fan-shaped tail in p.a. 225^{o}. J. Young reports that CCD frames taken with the Table Mountain 0.6-m reflector on Sept. 28.28–28.38 show a 4′′ coma without any central condensation and a very wide fan-shaped tail ∼ 6′′–8′′ long spanning p.a. 205^{o}–280^{o}, and a narrow 25′′ ‘spike’ in p.a. 255^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Sept.26.25518  3^{h}11^{m}20.^{s}77  +9^{o}32′37′′.4  18.3

Additional astrometry, preliminary parabolic orbital elements (T = 2004 Dec. 9.1 TT, ω = 92^{o}.5, Ω = 19^{o}.3, i = 114^{o}.6, equinox 2000.0, q = 0.687 AU), and an ephemeris appear on MPEC 2004-S81.

SUPERNOVAE 2004er AND 2004es Further to IAUC 8406, H. Pugh, S. Park, and W. Li report the LOSS

discoveries of two apparent supernovae on unfiltered KAIT images.

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004er Sept. 25.40 2^{h}36^{m}59.^{s}04 − 5^{o}21′19′′.7 17.0 7′′.3 E, 22′′.2 S
2004es Sept. 26.54 7 23 34.62 +41 25 38.3 17.6 15′′.9 E, 27′′.0 S

Additional approximate magnitudes: SN 2004er in MCG –01-7-24, Sept. 21.38 UT, [19.0; 26.38, 16.8. SN 2004es in UGC 3825, Apr. 8.21, [19.0; Sept. 27.53, 17.5. SN 1998C appeared 26′′ to the north of 2004es.

2004 September 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green